# Vector-Quantized Vision Foundation Models for Object-Centric Learning

### v5

@@ 4.2. Evaluate on Object Discovery
(ARI), ARIfg (foreground), mean Intersection-over-Union [^12] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.jaccard_score.html (mIoU) and mean Best Overlap (Uijlings et al., 2013) (mBO) as metrics.
@@
As shown in Table 3, whether using higher input resolution or different aggregators, VVO maintains its superiority over baselines. As shown in Table 4, VVO outperforms recent OCL general improvers, i.e., GDR and MSF. VVO also surpasses state-of-the-arts, i.e., SPOT (Kakogeorgiou et al., 2024) and VideoSAUR (Zadaianchuk et al., 2024), with their special types of decoding.
@@ 4.3. Evaluate on Set Prediction
Set prediction task directly shows how much object information those slots grasp. We use OCL to represent dataset COCO as slots, and use a small MLP to predict the object class label and bounding box corresponding to each slot by following this work (Seitzer et al., 2023). We measure top1 accuracy [^13] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.accuracy_score.ht
@@ 4.4. Deploy to the Downstream
Better object representation benefits downstream tasks. We follow the convention to pretrain OCL models on Physion and represent this dataset as slots. Then the object-centric dynamics model SlotFormer (Wu et al., 2023a) is trained on those slots in an auto-regressive manner along the time dimension. We use temporal versions of DINOSAUR and our VVO, i.e., VQDINO Mlp to extract slots.
@@ 5. Analysis
As shown in Figure 7, super-pixels in a feature map $\bm{Z}$ all belong to two objects $\bm{o}_{1}$ and $\bm{o}_{2}$, so two queries are needed for aggregation, which is basically sum of super-pixels weighted by normalized minus distances between queries and super-pixels (Locatello et al., 2020).

### v6

@@ 4.2. Evaluate on Object Discovery
(ARI), ARIfg (foreground), mean Intersection-over-Union [^12] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.jaccard_score.html (mIoU) and mean Best Overlap (uijlings2013selectivesearch) (mBO) as metrics.
@@
As shown in Table 3, whether using higher input resolution or different aggregators, VVO maintains its superiority over baselines. As shown in Table 4, VVO outperforms recent OCL general improvers, i.e., GDR and MSF. VVO also surpasses state-of-the-arts, i.e., SPOT (kakogeorgiou2024spot) and VideoSAUR (zadaianchuk2024videosaur), with their special types of decoding.
@@ 4.3. Evaluate on Set Prediction
Set prediction task directly shows how much object information those slots grasp. We use OCL to represent dataset COCO as slots, and use a small MLP to predict the object class label and bounding box corresponding to each slot by following this work (seitzer2023dinosaur). We measure top1 accuracy [^13] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.accuracy_score.ht
@@ 4.4. Deploy to the Downstream
Better object representation benefits downstream tasks. We follow the convention to pretrain OCL models on Physion and represent this dataset as slots. Then the object-centric dynamics model SlotFormer (wu2022slotformer) is trained on those slots in an auto-regressive manner along the time dimension. We use temporal versions of DINOSAUR and our VVO, i.e., VQDINO Mlp to extract slots.
@@ 5. Analysis
As shown in Figure 7, super-pixels in a feature map $\bm{Z}$ all belong to two objects $\bm{o}_{1}$ and $\bm{o}_{2}$, so two queries are needed for aggregation, which is basically sum of super-pixels weighted by normalized minus distances between queries and super-pixels (locatello2020slotattent).